OpenStudy (anonymous):

dd 25 mL of 20% Glucose Solution and 75 mL of Water to the third Beaker. How can you tell how many parts water?

OpenStudy (anonymous):

*add

OpenStudy (anonymous):

@amoodarya Help Me Please.

OpenStudy (anonymous):

??

OpenStudy (amoodarya):

i think 25(80 - 20)+75 water because 25*20% is glucose

OpenStudy (amoodarya):

sorry 25(100 - 20)+75 is water=20 +75=95ml water

OpenStudy (amoodarya):

i had a typing mistake first time i wrote 80 instead of 100

OpenStudy (anonymous):

where did you get 100 from.

OpenStudy (amoodarya):